# Conformal Survival Bands for Risk Screening under Right-Censoring

### v1

@@ 1.2 Preview of Our Contributions
We develop a method to construct principled and interpretable “uncertainty bands” around individual survival curves estimated by black-box models, using right-censored data. These bands, which we name conformal survival bands, are designed to support clinical screening tasks and are rigorously calibrated in a useful predictive sense, as explained below.
@@
Figure 1: Illustration of the use of conformal survival bands (shaded regions) for screening test patients in simulated censored data. Solid black curves show survival estimates from either an inaccurate (top) or accurate (bottom) survival forest model. Dashed green curves represent the true survival probabilities. The goal is to identify low-risk patients—those with more than 80% probability of surviving beyond 3 years (marked by the vertical line). A patient is flagged by our method if their entire conformal band lies above the 80% threshold (horizontal line) at this time, ensuring at least 80% of flagged individuals survive longer. Flagging decisions are indicated by colored markers: red triangles (our method), green asterisks (oracle), black circles (estimated model), and blue diamonds (Kaplan-Meier). Patients 2–4 are mistakenly flagged as low risk by the inaccurate model. With the accurate model, our method can identify truly low-risk patients (e.g., Patient 1), unlike the KM curve.
@@
Rather than attempting to directly estimate population-level parameters, our bands are calibrated for distribution-free predictive screening. They allow practitioners to identify individuals whose estimated survival probability lies below (or above) a clinically meaningful threshold—while guaranteeing that, among all such flagged individuals, the false discovery rate is approximately controlled in a precise predictive sense. For example, in Figure 1, we illustrate a scenario in which the practitioner wishes to identify “low-risk” patients—defined as those with more than an $p=80\%$ chance of surviving beyond time $t=5$ months. A patient is flagged if their conformal band lies entirely above the 80% line at $t=5$. Our method asymptotically guarantees that, among all flagged individuals, on average at least 80% will survive beyond this time point. This guarantee holds pointwise over all fixed choices of the survival probability threshold $p$ and time horizon $t$, and for “low-risk” as well as “high-risk” screening, making the method flexible and broadly applicable.
@@
This predictive notion of uncertainty is well-aligned with how survival curves are commonly used in real-world settings: to guide actionable decisions such as prioritizing patients for further testing or early intervention. However, black-box survival models typically provide only point estimates, without any reliable measure of uncertainty or formal calibration guarantees. As in the example shown, screening based on point predictions alone, or even KM estimates, may lead to too many patients incorrectly labeled as low risk—highlighting the practical benefits of our calibrated bands, which offer rigorous control over such errors.
Our approach integrates three key concepts: conformal inference (Vovk et al., 2005; Fontana et al., 2023), inverse probability of censoring weighting (IPCW) (Robins and Rotnitzky, 1992; Robins, 1993), and false discovery rate (FDR) control (Benjamini and Hochberg, 1995). Concretely, we build on tests for random null hypotheses of the type:
$H_{\mathrm{lt}}(t):T\geq t,\quad\text{or}\quad H_{\mathrm{rt}}(t):T\leq t,$
@@
where $T$ is the (unobserved) survival time of the test individual and $t>0$ is a user-specified time threshold. In the following, we will compute conformal p-values to test these hypotheses and show how to use them to construct the conformal survival bands previewed in Figure 1. Although each of the statistical concepts upon which we build already exists on its own, the way we integrate them is original and supported by novel theoretical results.
@@ 2.4 Conformal Survival Bands for Screening High- or Low-Risk Individuals
3:  Estimate censoring weights $\hat{w}(t,x)$ using ^$\hat{\mathcal{M}}_{C}$ as in Eq. (6).
@@ 2.5 Theoretical Guarantees for Calibrated Survival Bands
An analogous result holds for the right-tail p-values $\hat{\phi}_{\mathrm{rt}}(t;X_{n+j})$, used to test $H_{0,j}^{\mathrm{rt}}$.
@@
2.6 Extension: Doubly Robust Survival Bands
@@
We now present a simple but practically valuable extension that can improve robustness to misspecification of the censoring model. Specifically, this variant aims for double robustness: it yields approximately calibrated screening rules as long as either the censoring model or the survival model is consistent—though not necessarily both. While we do not pursue a formal proof of double robustness here, the underlying idea is intuitive.
@@ 3.1 Setup
The first three settings are borrowed from Sesia and Svetnik (2024) and are ordered by decreasing modeling difficulty. The first two simulate challenging scenarios where the true survival distribution is highly nonlinear or complex, making accurate model fitting difficult and highlighting the need for robust uncertainty estimation via conformal inference. The third setting, originally appearing in Candès et al. (2023), corresponds to a simpler survival distribution that is easier to learn with standard survival models, reducing the gap between model-based and conformal prediction. Finally, the fourth setting, which is new, introduces a covariate shift between the training and calibration/test distributions, leading to potential model misspecification even if the survival model fits the training distribution well, and further emphasizing the value of conformal inference methods such as CSB.
@@
Our goal is to evaluate the proposed conformal survival band (CSB) method and compare it to natural baseline approaches for screening test patients according to rules of the form $P[T>t]>p$ (for low-risk selection) or $P[T>t]<p$ (for high-risk selection), evaluated at various time thresholds $t$ and probability levels $p$. We consider three methods for screening patients based on estimated conditional survival probabilities: (i) model, using the point estimates of conditional survival probabilities computed by a black-box survival model ^$\hat{\mathcal{M}}^{\text{surv}}$ fitted on all available data (training and calibration); (ii) KM, using a Kaplan-Meier estimator fitted on the calibration data; and (iii) CSB, using the conformal survival band produced by Algorithm 1, implemented using the doubly robust extension described in Section 2.6. In addition, we compare against an ideal oracle that uses the true conditional survival probabilities from the data-generating distribution; although not practical, the oracle provides valuable insight into the achievable performance limits.
@@
All experiments are repeated 100 times, and results are averaged across repetitions.
@@
We consider four model families for fitting the censoring and survival models, denoted respectively by ^$\hat{\mathcal{M}}^{\text{cens}}$ and ^$\hat{\mathcal{M}}^{\text{surv}}$, to ensure consistent comparisons across different calibration methods. The model families are: (1) grf, a generalized random forest (R package grf); (2) survreg, an accelerated failure time (AFT) model with a log-normal distribution (R package survival); (3) rf, a random survival forest (R package randomForestSRC); and (4) cox, the Cox proportional hazards model (R package survival). The survival model ^$\hat{\mathcal{M}}^{\text{surv}}$ is used both by the model screening method and as input to the proposed conformal survival band (CSB) method. In contrast, the censoring model ^$\hat{\mathcal{M}}^{\text{cens}}$ is only used within the CSB method, to compute the weights for constructing conformal survival bands.
@@ Appendix A Mathematical Proofs
and the p-values $p_{j}^{*}$ are mutually independent (conditional on the calibration data). Therefore, by Theorem A.14, the BH procedure applied to $(p_{1}^{*},\ldots,p_{m}^{*})$ controls the FDR at level $\alpha$:
@@ A.3.4 Stability and validity of the BH procedure
(Jin and Candès, 2023b, Thm. 2.3) Let $(H_{1},\dots,H_{m})\in\{0,1\}^{m}$ be random variables indicating the status of $m$ hypotheses, where $H_{j}=1$ means that the $j$th hypothesis is a true null. Let $(p_{1},\dots,p_{m})\in[0,1]^{m}$ be random variables, interpreted as p-values for the corresponding hypotheses. Assume:
@@ B.1 Additional Details on Data Generating Distributions
Table 3: Summary of four synthetic data generation settings considered for numerical experiments. Settings 1, 2, and 3 are adapted from Sesia and Svetnik (2024), with Setting 3 originally appearing in Candès et al. (2023). Setting 4 is new.
@@ B.6 Effect of Different Censoring and Survival Models
Table 4: Performance of different methods for screening low-risk patients with $P(T>6)>0.80$ in a challenging synthetic data scenario (Setting 1 from Table 3), using various censoring and survival models. The training sample size is fixed at 1000. Other details are as in Figure 2.
@@ Appendix C Additional Details on Real Data Applications
Table 13: Detailed screening results for high-risk selection using the Cox model, with threshold rule $P(T>t_{1})<0.80$ and $t_{1}$ set to the 0.1 quantile of observed times in each dataset. Shown are the screened proportion and bounds on the survival rate among selected patients, aggregated over 100 repetitions.

### v2

@@ 1.2 Preview of Our Contributions
We develop a method to construct principled and interpretable “uncertainty bands” around individual survival curves estimated by black-box models, using right-censored data. These bands, which we call conformal survival bands, are designed to support clinical screening tasks within a multiple testing framework and are rigorously calibrated in a useful predictive sense, as explained below.
Here, “multiple testing” refers to the fact that our method constructs uncertainty bands for each of several individuals in a test set by processing the entire test set jointly. This is analogous to how the Benjamini–Hochberg procedure processes all $p$-values for multiple hypotheses at once to control the false discovery rate (FDR) (Benjamini and Hochberg, 1995). The multiple testing nature of our approach will become clearer once we present its details and the theoretical guarantees it provides.
@@
Figure 1: Illustration of the use of conformal survival bands (shaded regions) for screening test patients in simulated censored data. Solid black curves show survival estimates from either an inaccurate (top) or accurate (bottom) survival forest model. Dashed green curves represent the true survival probabilities. The goal is to identify low-risk patients—those with more than 80% probability of surviving beyond 3 years (vertical line). A patient is flagged by our method if their entire conformal band lies above the 80% threshold (horizontal line) at this time, ensuring at least 80% of flagged individuals survive longer. Flagging decisions are indicated by colored markers: red triangles (our method), green asterisks (oracle), and black circles (estimated model). Kaplan-Meier makes no selections here. Patients 2–4 are mistakenly flagged as low risk by the inaccurate model. With the accurate model, our method can identify truly low-risk patients (e.g., Patient 1), unlike the KM curve.
@@
Rather than attempting to directly estimate population-level parameters, our bands are calibrated for distribution-free predictive screening. They allow practitioners to identify individuals whose estimated survival probability lies below or above a clinically meaningful threshold, while guaranteeing that among all such flagged individuals, the false discovery rate is approximately controlled in a precise predictive sense. This is why our method is intrinsically a multiple testing approach that processes the entire test set jointly.
For example, in Figure 1, we illustrate a scenario in which the practitioner wishes to identify “low-risk” patients—defined as those with more than an $p=80\%$ chance of surviving beyond time $t=3$ years. A patient is flagged if their conformal band lies entirely above the 80% line at $t=3$. Our method asymptotically guarantees that, among all flagged individuals, on average at least 80% will survive beyond this time point. This guarantee holds pointwise over all fixed choices of the survival probability threshold $p$ and time horizon $t$, and for “low-risk” as well as “high-risk” screening, making the method flexible and broadly applicable.
@@
This predictive notion of uncertainty within a multiple testing framework is well-aligned with how survival curves are commonly used in real-world settings: to guide actionable decisions such as prioritizing patients for further testing or early intervention. However, black-box survival models typically provide only point estimates, without any reliable measure of uncertainty or formal calibration guarantees. As in the example shown, screening based on point predictions alone, or even KM estimates, may lead to too many patients incorrectly labeled as low risk—highlighting the practical benefits of our calibrated bands, which offer rigorous control over such errors.
@@
Our approach integrates three key concepts: conformal inference (Vovk et al., 2005; Fontana et al., 2023), inverse probability of censoring weighting (IPCW) (Robins and Rotnitzky, 1992; Robins, 1993), and FDR control (Benjamini and Hochberg, 1995). Concretely, we build on tests for random null hypotheses stating that $T\geq t$ or $T\leq t$, where $T$ is the (unobserved) survival time of the test individual and $t>0$ is a user-specified time threshold. We will compute conformal p-values to test these hypotheses and show how to use them to construct the conformal survival bands previewed in Figure 1. Although each of the statistical concepts upon which we build already exists on its own, the way we integrate them is original and supported by novel theoretical results.
@@ 2.4 Conformal Survival Bands for Screening High- or Low-Risk Individuals
3:  Estimate censoring weights $\hat{w}(t,x)$ for the data in $\mathcal{D}_{\text{cal}}^{\prime}$ using ^$\hat{\mathcal{M}}_{C}$ as in Eq. (6).
@@ 2.5 Theoretical Guarantees for Calibrated Survival Bands
An analogous result holds for the right-tail p-values $\hat{\phi}_{\mathrm{rt}}(t;X_{n+j})$, used to test $H_{\mathrm{rt}}(t;j)$.
@@
2.6 Extension: Making the Survival Bands Doubly Robust
@@
We now describe a simple extension that makes our method more conservative, aiming to improve robustness to misspecification of the censoring model. Specifically, this variant targets double robustness: it is designed to deliver approximately calibrated screening rules as long as either the censoring model or the survival model is consistent—though not necessarily both. We do not prove a formal double robustness guarantee, as the paper is already quite long. Nonetheless, we believe that a formal double robustness result for this extension could be established using techniques similar to those in Sesia and Svetnik (2025).
@@ 3.1 Setup
The first three settings are borrowed from Sesia and Svetnik (2025) and are ordered by decreasing modeling difficulty. The first two simulate challenging scenarios where the true survival distribution is highly nonlinear or complex, making accurate model fitting difficult and highlighting the need for robust uncertainty estimation via conformal inference. The third setting, originally appearing in Candès et al. (2023), corresponds to a simpler survival distribution that is easier to learn with standard survival models, reducing the gap between model-based and conformal prediction. Finally, the fourth setting, which is new, introduces a covariate shift between the training and calibration/test distributions, leading to potential model misspecification even if the survival model fits the training distribution well, and further emphasizing the value of conformal inference.
@@
Our goal is to evaluate the proposed conformal survival band (CSB) method and compare it to natural baseline approaches for screening test patients according to rules of the form $P[T>t]>p$ (for low-risk selection) or $P[T>t]<p$ (for high-risk selection), evaluated at various time thresholds $t$ and probability levels $p$. We consider three methods for screening patients based on estimated conditional survival probabilities: (i) model, using the point estimates of conditional survival probabilities computed by a black-box survival model ^$\hat{\mathcal{M}}^{\text{surv}}$ fitted on all available data (training and calibration); (ii) KM, using a Kaplan–Meier estimator fitted on the calibration data; and (iii) CSB, using the conformal survival band produced by Algorithm 1, implemented with the doubly robust extension described in Section 2.6.
We focus on the doubly robust implementation of CSB in our experiments because it is more interpretable, ensuring that the predicted survival probability always lies within the band, as a practitioner would likely expect. Although formally studying the double robustness property of this extension is beyond the scope of this paper, this approach is already supported by the statistical guarantees established in Section 2.5, since it is always more conservative than the baseline method described in Section 2.4.
In addition, we compare against an ideal oracle that uses the true conditional survival probabilities from the data-generating distribution; although not practical, the oracle provides valuable insight into the achievable performance limits.
@@
All experiments are independently repeated 100 times, and results are reported as means across repetitions, with error bars representing two standard errors (SE).
@@
We consider four model families for fitting the censoring and survival models, denoted respectively by ^$\hat{\mathcal{M}}^{\text{cens}}$ and ^$\hat{\mathcal{M}}^{\text{surv}}$, to ensure consistent comparisons across different calibration methods. The model families are: (1) grf, a generalized random forest (R package grf); (2) survreg, an accelerated failure time (AFT) model with a log-normal distribution (R package survival); (3) rf, a random survival forest (R package randomForestSRC); and (4) cox, the Cox proportional hazards model (R package survival). The survival model ^$\hat{\mathcal{M}}^{\text{surv}}$ is used both by the model screening method and as input to the CSB method. In contrast, the censoring model ^$\hat{\mathcal{M}}^{\text{cens}}$ is only used within the CSB method, to compute the weights for constructing conformal survival bands.
@@ 5 Discussion
An additional limitation is that our method can be conservative in some settings, resulting in fewer selections than the ideal oracle or even fewer than the black-box survival model when that model appears well calibrated. A possible way to mitigate this over-conservativeness in future work would be to incorporate an adaptive estimator of the proportion of null hypotheses (Gao, 2024), similar to how Storey’s method extends the BH procedure (Storey et al., 2004).
@@
Acknowledgements
We thank the two anonymous referees for their detailed feedback and valuable suggestions on an earlier version of this manuscript.
@@ Appendix A Mathematical Proofs
and the p-values $p_{j}^{*}$ are mutually independent (conditional on the calibration data). Therefore, by Theorem A.14, BH applied to $(p_{1}^{*},\ldots,p_{m}^{*})$ controls the FDR below $\alpha$:
@@ A.3.4 Stability and validity of the BH procedure
(Jin and Candès, 2023b, Thm. 3) Let $(H_{1},\dots,H_{m})\in\{0,1\}^{m}$ be random variables indicating the status of $m$ hypotheses, where $H_{j}=1$ means that the $j$th hypothesis is a true null. Let $(p_{1},\dots,p_{m})\in[0,1]^{m}$ be random variables, interpreted as p-values for the corresponding hypotheses. Assume:
@@ B.1 Additional Details on Data Generating Distributions
Table 3: Summary of the four synthetic data generation settings considered for numerical experiments. Settings 1, 2, and 3 are adapted from Sesia and Svetnik (2025), with Setting 3 originally introduced in Candès et al. (2023). Setting 4 is new. In this table, $X$ denotes a $p$-dimensional vector of covariates for a random individual, and $X_{j}$ indicates its $j$-th element for any $j\in[p]$. Not all covariates are used to generate $Y$; most serve as noisy variables, making it more challenging to fit accurate survival and censoring models. In some experiments, we simplify the training process by excluding a subset of these noisy variables.
@@ B.6 Effect of Different Censoring and Survival Models
Table 4: Performance of different methods for screening low-risk patients with $P(T>6)>0.80$ in a challenging synthetic scenario (Setting 1 from Table 3), using various censoring and survival models. The training sample size is 1000. Other details are as in Figure 2.
@@ Appendix C Additional Details on Real Data Applications
Table 13: Detailed screening results for high-risk selection using the Cox model, with threshold rule $P(T>t_{1})<0.80$ and $t_{1}$ set to the 0.1 quantile of observed times in each dataset. Shown are the screened proportion and bounds on the survival rate among selected patients, aggregated over 100 repetitions. In some cases, no patients are selected.